OpenStudy (anonymous):

simplify square root of 512k^4

OpenStudy (istim):

Do you use logarithmic here?

OpenStudy (anonymous):

logawhat?

OpenStudy (istim):

No?

OpenStudy (istim):

Ok...

OpenStudy (anonymous):

so can u help me?

OpenStudy (istim):

Probably not...

OpenStudy (campbell_st):

\[\sqrt{(16k^4 x 2) } = 4k^2\sqrt{2}\]

OpenStudy (anonymous):

how do you do that?

OpenStudy (campbell_st):

oops \[\sqrt{512k^4} = \sqrt{256k^4 \times2} = 15k^2\sqrt{?}\]

OpenStudy (campbell_st):

? = 2

OpenStudy (anonymous):

2

OpenStudy (campbell_st):

and 16 not 15

OpenStudy (anonymous):

so what's the answer?

OpenStudy (campbell_st):

\[16k^2\sqrt{2}\]